# COMMUNITYNOTES: A Dataset for Exploring the Helpfulness of Fact-Checking Explanations

### v1

@@ Abstract
Fact-checking on major platforms, such as X, Meta, and TikTok, is shifting from expert-driven verification to a community-based setup, where users contribute explanatory notes to clarify why a post might be misleading. An important challenge here is determining whether an explanation is helpful for understanding real-world claims and the reasons why, which remains largely underexplored in prior research. In practice, most community notes remain unpublished due to slow community annotation, and the reasons for helpfulness lack clear definitions. To bridge these gaps, we introduce the task of predicting both the helpfulness of explanatory notes and the reason for this. We present COMMUNITYNOTES, a large-scale multilingual dataset of 104k posts with user-provided notes and helpfulness labels. We further propose a framework that automatically generates and improves reason definitions via automatic prompt optimization, and integrate them into prediction. Our experiments show that the optimized definitions can improve both helpfulness and reason prediction. Finally, we show that the helpfulness information are beneficial for existing fact-checking systems. The code and data are available at: https://github.com/ruixing76/Helpfulness-FCExp.
COMMUNITYNOTES: A Dataset for Exploring the Helpfulness of Fact-Checking Explanations
@@
Figure 1: An example of X Community Notes. The user-generated note appears as “Readers added context.”
@@ 1 Introduction
Beyond simple verification, fact-checking systems should provide explanations that clarify why a claim is misleading (Warren et al., 2025). Explanations not only strengthen user trust, but also mitigate pitfalls of fact-checking, e.g., backfire, where the false belief is reinforced rather than hindered (Lewandowsky et al., 2012; Guo et al., 2022; Xing et al., 2025). Typically, fact-checking explanations are produced by expert organizations. Yet the ecosystem is currently shifting, and platforms are increasingly experimenting with community-based fact-checking, where contributors collaboratively provide explanatory notes on claims (Augenstein et al., 2025). This crowdsourcing process distributes the labor of verification and offers diverse perspectives from the community. A representative example is Community Notes on X (formerly Twitter), which allows users to attach notes explaining potentially misleading posts (Figure˜1). [^1] https://communitynotes.x.com/guide/
@@
Critically, for a note to become visible or formally published, it must receive sufficient amount of ratings from a diverse set of users (De et al., 2025). A key challenge in that respect is to decide whether an explanation is helpful for understanding real-world claims and why, which is underexplored in previous research on fact-checking explanations (Atanasova et al., 2020; Guo et al., 2022; Russo et al., 2023; Eldifrawi et al., 2024). This process is also slow on Community Notes: notes can take from hours to days to appear, and the vast majority (over 90%) never become visible, despite the invested contributors efforts. Moreover, while the platform provides reasons for judging notes as helpful or unhelpful, these reasons lack clear definitions, leaving both raters and contributors uncertain about the evaluation criteria (De et al., 2025; Augenstein et al., 2025). This opacity hinders the efficiency, scalability, and explainability of crowd-sourced fact-checking.
Motivated by the above, here we explore the idea of simplifying and making more explicit the helpfulness annotation process. Our contributions are as follows:
@@
  • We introduce the task of predicting the helpfulness and its reasons given a note and the original X post.
@@
  • We present COMMUNITYNOTES, a large-scale explanation helpfulness dataset containing over 104K potentially misleading posts and corresponding user-provided notes.
@@
  • We propose a framework that automatically generates and optimizes reason label definitions via prompt optimization; integrating them to helpfulness/reason predictors to improve their performance.
@@
  • We apply helpfulness prediction to the evidence sufficiency task (Atanasova et al., 2022) and further incorporate it as auxiliary signals into automated fact-checking on CLIMATE-FEVER (Diggelmann et al., 2020), showing its potential to enhance existing fact-checking systems.
@@ 2 Related Work
The proliferation of misinformation has become a growing societal concern, posing significant risks to areas such as public health, science, and politics (Borenstein et al., 2025). To address this challenge, fact-checking has emerged as a central strategy in mitigating the spread of false or misleading information (Guo et al., 2022).
Traditionally, fact-checking has been conducted by professional third-party organizations that evaluate the veracity of claims through expert analysis. However, recent years have witnessed a paradigm shift toward more decentralized, community-based approaches (Renault et al., 2024; Augenstein et al., 2025). Major social media platforms have increasingly embraced this model. X’s Community Notes has become the most prominent example, enabling users to collaboratively annotate and assess the accuracy of online content. Following this trend, Meta recently terminated its partnerships with professional fact-checkers in favor of a similar community-driven system (Catalanello and Sanders, 2025), while TikTok has also introduced a comparable feature (TikTok, 2025).
Empirical evidence suggests that community-based fact-checking can be highly effective in countering misinformation. Studies have shown that the presence of a Community Note can reduce the rate of retweets by nearly half (Renault et al., 2024), while displaying such notes on misleading posts has been found to decrease their spread by an average of 61% (Chuai et al., 2024). Moreover, authors of posts flagged by community notes are up to 80% more likely to delete their content, and such posts generally exhibit markedly lower virality (Drolsbach and Pröllochs, 2023; Renault et al., 2024). However, Community Notes often requires a long time to gather sufficient contributor ratings to determine helpfulness, and the absence of reason definitions further undermines the trustworthiness of the mechanism.
@@ 3.1 Task Definition
Given a post $P$ that contains potentially misleading information and a note $N$ that provides an explanation about that post, the task is to predict two types of labels: a helpfulness label $L_{helpful}$, indicating whether the note is helpful in explaining the misleading nature of the post, and a reason label $L_{reason}$, specifying why the note is (un)helpful. The helpfulness prediction is formulated as a binary classification task with the label space $\{Helpful,Unhelpful\}$. The reason prediction is a multi-label classification task, where the predefined label set includes eight reasons for helpful notes and ten reasons for unhelpful ones (See Table˜14 for detailed reason labels).
3.2 Data Collection and Pre-processing
We collected all our data from official Community Notes website between January 2021 and December 2024.222https://communitynotes.x.com/guide/en/under-the-hood/download-data We construct our dataset by combining all publicly available data releases, which include the Notes, Ratings, Note Status History, and User Enrollment tables. Since the ratings data are provided in multiple shards due to their large volume, we first merge all shards and join them with the other components to obtain complete note–rating pairs. We then apply the official note-ranking algorithm to compute the aggregated helpfulness and reason labels for each note. [^3] https://communitynotes.x.com/guide/en/under-the-hood/ranking-notes For helpfulness prediction, we remove entries labeled as NEED_MORE_RATINGS, leaving the helpfulness label space binary—CURRENTLY_RATED_HELPFUL or CURRENTLY_RATED_NOT_HELPFUL. The reason annotation is multi-label, covering 18 predefined categories. Finally, we retrieve tweet content and metadata via the X API and web crawling (Borenstein et al., 2025), linking each post with its corresponding notes to form the final dataset used in our experiments.
@@ 3.2 Data Collection and Pre-processing
Table 1: Data split statistics for English and other language subsets (updated).
Table˜1 provides the data distribution. Our final dataset contains 104,966 posts, including 58,569 in English and 46,397 in other languages. We perform stratified partitioning with a 7:1:2 ratio for training, development, and testing sets, maintaining separate splits for English and other language data. Henceforth we will refer to this dataset as COMMUNITYNOTES.
3.3 Dataset Statistics
@@
We inspected multiple aspects of our entire COMMUNITYNOTES dataset. Table˜2 shows the token length distribution for posts and notes correspondingly. The median token number are 42 for posts and 70 for notes, which suggests most posts and notes are short text snippets.
@@
Mixed helpful / unhelpful notes 4,557 4.34%
@@
In Table˜3, we observe that 84.74% of posts have only one note, 11.84% have two notes, and the remaining posts contain three or more notes. Table˜4 shows that 64.2% of posts have notes that are all rated as helpful, 31.4% have only unhelpful notes, and the remaining 4.4% include notes with mixed ratings. There are eight predefined reasons for a note being rated as helpful and ten reasons for being rated as unhelpful. The most common helpful reasons are helpfulAddressClaim, helpfulImportantContext, and helpfulClear, whereas nothelpfulArgumentativeOrBiased and nothelpfulMissingKeyPoints are the most frequent unhelpful reasons (See detailed reason distribution in Appendix Figure˜4).
Figure˜2 illustrates the language distribution of COMMUNITYNOTES. English is the dominant language, accounting for 57.3% of all notes, followed by Japanese (10.5%) and Spanish (9.9%). This indicates that our final dataset effectively captures the multilingual nature of crowd-sourced explanations.
Figure 2: The languages in the COMMUNITYNOTES: fr-French, es-Spanish, ja-Japanese, pt-Portuguese, de-German, Other: note languages that appear less than 1,000 times.
Beyond analyzing note statistics, we also examine the popularity of the posts targeted by these notes. We use the number of replies, likes, and retweets as metrics (See detailed distributions in Appendix Figure˜5). These posts receive a median of approximately 3k likes, 300 replies, and 700 retweets. This indicates that the posts attracting community notes are highly popular. Given that they are flagged as potentially misleading, this underscores the urgent need to provide helpful explanations for their claims.
@@ 3.4 Benchmarking
To establish a baseline for understanding the model performance in predicting the helpfulness of the notes, we conducted a comprehensive set of experiments using both Small Language Models (SLMs) (BERT-style encoder-only architectures in our paper) and decoder-only Large Language Models (LLMs). For SLMs, we formulate the task as multi-task classification, where the model is jointly trained to predict (1) a binary helpfulness label and (2) a multi-label reason classification across 18 predefined categories. The model input is constructed as a concatenated sequence in the format “[CLS] Claim: {claim} [SEP] Note: {note_text}” and subsequently passed through two separate classification heads. For LLMs, we fine-tune models in the same multi-task way (See detailed prompt in Appendix Table˜14). We partitioned our experiments by model and language setting. For English, we experimented with SLMs including BERT, ModernBERT, RoBERTa and DeBERTa (both base and large). We also evaluated the performance of LLMs including Llama3.1-8B and Mistral-7B, under both zero-shot and LoRA fine-tuning conditions (Hu et al., 2022). For the multilingual part, we used models specifically trained on multilingual data, including BERT-base-multilingual (mBERT) and XLM-RoBERTa (base and large).
Language Model Helpfulness F1 Reason P Reasons R Reason F1
@@
Llama3.1-8B-ins 0.928 0.653 0.652 0.653
Mistral-7B-ins 0.926 0.647 0.647 0.647
Table 5: Benchmark results for helpfulness and reasons prediction. Helpfulness means the binary classification of notes helpfulness. Reason means multi-label classification of helpfulness reasons.
@@
The models generally perform well on the helpfulness prediction task, with most of them achieving F1 scores beyond 0.88. This suggests that distinguishing between helpful and non-helpful notes is a relatively easy task. In contrast, the reasons prediction task is much more challenging. This is showed by the much lower F1 scores, particularly the F1 scores which consistently fall below 0.70.
Comparing SLMs to LLMs: The table shows that LLMs like Llama3.1-8B and Mistral-7B outperform smaller SLMs on Helpfulness prediction. The Mistral-7B model achieves the highest F1 Score (0.920), demonstrating the superior capabilities of larger, more advanced models. However, the LLMs perform worse than SLMs in reason prediction, with their F1 scores being comparable to the best-performing SLMs like DeBERTa-large. Within the SLM families (BERT, RoBERTa, DeBERTa), larger models consistently outperform their base-sized counterparts. For example, RoBERTa-large achieves better scores than RoBERTa-base across all metrics. This highlights the importance of model capacity for capturing the nuances required for this complex task. The table also shows that DeBERTa models generally perform at or near the top among the SLMs, particularly for the reasons task with the highest F1 Score of 0.665 among all SLMs.
@@
Figure 3: Our framework for automatic reason definition generation and optimization.

### v2

@@ Abstract
Fact-checking on major platforms, such as X, Meta, and TikTok, is shifting from expert-driven verification to a community-based setup, where users contribute explanatory notes to clarify why a post might be misleading. An important challenge here is determining whether an explanation is helpful for understanding real-world claims and the reasons why, which remains largely underexplored in prior research. In practice, most community notes remain unpublished due to slow community annotation, and the reasons for helpfulness lack clear definitions. To bridge these gaps, we introduce the task of predicting both the helpfulness of explanatory notes and the reason for this. We present COMMUNITYNOTES, a large-scale multilingual dataset of 104k posts with user-provided notes and helpfulness labels. We further propose a framework that automatically generates and improves reason definitions via automatic prompt optimization, and integrate them into prediction. Our experiments show that the optimized definitions can improve both helpfulness and reason prediction. Finally, we show that the helpfulness information is beneficial for existing fact-checking systems. The code and the data are available at https://github.com/ruixing76/Helpfulness-FCExp.
COMMUNITYNOTES:
A Dataset for Exploring the Helpfulness of Fact-Checking Explanations
@@
Figure 1: An example of X Community Notes. The potential misleading tweet states, “mRNA Vaccines Permanently Fuse Into Human DNA”. The user-generated note appears under Readers added context and provides an explanation for the tweet.
@@ 1 Introduction
Beyond simple verification, fact-checking systems should provide explanations that clarify why a claim is misleading (Atanasova et al., 2020; Xing et al., 2022; Russo et al., 2023; Warren et al., 2025). Explanations not only strengthen user trust in fact-checking decisions, but also help mitigate common pitfalls of fact-checking, e.g., backfire effect, where exposure to corrections can paradoxically reinforce false beliefs rather than hinder them (Lewandowsky et al., 2012; Guo et al., 2022; Xing et al., 2025).
@@
Typically, fact-checking explanations are produced by expert organizations. Yet the ecosystem is currently shifting, and platforms are increasingly experimenting with community-based fact-checking, where contributors collaboratively provide explanatory notes on claims (Augenstein et al., 2025). This crowd-sourcing process distributes the labor of verification and offers diverse perspectives from the community. A representative example is Community Notes [^1] https://communitynotes.x.com/guide/ on X (formerly Twitter), which allows users to attach notes to explain potentially misleading posts. As illustrated in Figure˜1, the user-generated note appears under Readers added context and provides an explanation for the tweet.
Critically, for a note to become visible or formally published, it must receive a sufficient number of ratings from a diverse set of users (De et al., 2025). A key challenge in this process is determining whether an explanation is helpful for understanding real-world claims, as well as identifying why it is helpful, a question that remains largely underexplored in prior research on fact-checking explanations (Atanasova et al., 2020; Guo et al., 2022; Russo et al., 2023; Eldifrawi et al., 2024). In practice, this lack of clarity makes it difficult for contributors to write effective notes and for raters to assess them consistently.
@@
This process is slow on Community Notes: notes can take hours or even days to appear, and the vast majority (over 90%) never become visible, despite contributor efforts (De et al., 2025). Moreover, although the platform provides reasons for judging notes as helpful or unhelpful, these reasons lack clear definitions, leaving both contributors and raters unclear about the underlying evaluation criteria (De et al., 2025; Augenstein et al., 2025). This opacity hinders the efficiency, scalability, and explainability of the crowd-sourced fact-checking. Motivated by these challenges, we explore the idea of simplifying and making the helpfulness annotation process more explicit. Our contributions of this paper are as follows:
@@
  • We introduce the task of predicting the helpfulness of an explanatory note and the reasons underlying this judgment, given the note and the original X post.
@@
  • We present COMMUNITYNOTES, a large-scale dataset for explanation helpfulness, containing over 104K potentially misleading posts with corresponding user-provided notes.
@@
  • We propose a framework that automatically generates and optimizes reason label definitions via prompt optimization, and integrates these definitions into helpfulness and reason predictors to improve their performance.
@@
  • We apply explanation helpfulness prediction to the evidence sufficiency task (Atanasova et al., 2022) and further incorporate it as an auxiliary signal into automated fact-checking on CLIMATE-FEVER (Diggelmann et al., 2020), demonstrating its potential to enhance current existing automated fact-checking systems.
@@ 2 Related Work
The proliferation of misinformation has become a growing societal concern, posing significant risks to areas such as public health, science, and politics (Borenstein et al., 2025). Misleading claims can spread rapidly online, shaping public opinion and undermining trust in institutions. To address this challenge, fact-checking has emerged as a central strategy for mitigating the spread of false or misleading information, by evaluating claims and providing corrective context to users (Guo et al., 2022; Nakov et al., 2021; Xie et al., 2025).
Traditionally, fact-checking has been carried out by professional third-party organizations that assess claim veracity through expert analysis. However, as a result, recent years have seen a paradigm shift toward more decentralized, community-based approaches (Renault et al., 2024; Augenstein et al., 2025). Major social media platforms have increasingly adopted this paradigm. X’s Community Notes is the most prominent example, allowing their users to collaboratively annotate and assess online content by attaching contextual notes to potentially misleading posts. Following this trend, Meta ended its partnerships with professional fact-checkers [^2] Meta is ending its third-party fact-checking partnership with US partners. Here’s how that program works. in favor of a similar community-driven system, while TikTok has also introduced a comparable feature. [^3] Testing a New Feature to Enhance Content on TikTok.
Empirical evidence suggests that community-based fact-checking can be highly effective in countering misinformation at scale. Studies have shown that the presence of a Community Note can reduce the rate of retweets by nearly half (Renault et al., 2024), while displaying such notes on misleading posts has been found to decrease their spread by an average of 61% (Chuai et al., 2024). Moreover, authors of posts flagged by community notes are up to 80% more likely to delete their content, and such posts generally exhibit markedly lower virality (Drolsbach and Pröllochs, 2023; Renault et al., 2024). These findings highlight the potential of community-based explanations in debunking online misinformation.
Despite these benefits, Community Notes often require a long time to accumulate sufficient contributor ratings to determine helpfulness. In addition, the absence of reason definitions further undermines the trustworthiness of the mechanism.
@@ 3.1 Task Definition
Given a post $P$ that contains potentially misleading information and a note $N$ that provides an explanation about that post, the task is to predict two types of labels: a helpfulness label $L_{\mathit{helpful}}$, indicating whether the note is helpful in explaining the misleading nature of the post, and a reason label $L_{\mathit{reason}}$, specifying why the note is (un)helpful. The helpfulness prediction is formulated as a binary classification task with the label space $\{\mathit{Helpful},\mathit{Unhelpful}\}$. The reason prediction is a multi-label classification task, where the predefined label set includes eight reasons for helpful notes and ten reasons for unhelpful ones (see Appendix LABEL:tab:original_prompt for detailed reason labels).
@@
Table 1: COMMUNITYNOTES data split statistics for English and Other languages.
@@
3.2 Data Collection and Pre-processing
We collected all our data from X’s official Community Notes platform between January 2021 and December 2024, combining all publicly available releases. Because ratings are provided in multiple shards, we merge them and join them with the remaining components to obtain complete note–rating pairs. We then apply the official note-ranking algorithm to compute the aggregated helpfulness and reason labels. For helpfulness prediction, we remove entries labeled NEED_MORE_RATINGS, resulting in a binary label space–CURRENTLY_RATED_HELPFUL or CURRENTLY_RATED_NOT_HELPFUL. The reason annotation is multi-label, covering 18 predefined categories. Finally, we retrieve tweet content and metadata via the X API and web crawling (Borenstein et al., 2025), linking each post with its corresponding notes to create the final dataset. See Section˜A.2 for details of the preprocessing.
Table˜1 provides the data distribution. Our final dataset contains 104,966 posts, including 58,569 posts in English and 46,397 in other languages. Note that each post may be associated with one or more notes. We perform stratified partitioning with a ratio of 7:1:2 for training, development, and testing sets, maintaining separate splits for English and other language data. Henceforth we will refer to this dataset as COMMUNITYNOTES.
@@ 3.2 Data Collection and Pre-processing
Figure 2: The histogram of the Note Reason label distribution. COMMUNITYNOTES contains 18 reason categories—8 corresponding to helpful notes and 10 to not helpful notes.
3.3 Dataset Statistics
We inspected multiple aspects of our entire COMMUNITYNOTES dataset. Table˜2 shows the token length distribution for posts and notes correspondingly. The median token numbers are 42 for posts and 70 for notes, which suggests most posts and notes are short text snippets in our dataset.
In Table˜3, we observe that 84.74% of posts have only one note, 11.84% have two notes, and the remaining posts contain three or more notes. Table˜4 shows that 64.36% of posts have notes that are all rated as helpful, 31.30% have only unhelpful notes, and the remaining 4.34% include notes with mixed ratings. There are eight predefined reasons for a note being rated as helpful and ten reasons for being rated as unhelpful. The most common helpful reasons are helpfulAddressClaim, helpfulImportantContext, and helpfulClear, whereas nothelpfulArgumentativeOrBiased and nothelpfulMissingKeyPoints are the most frequent unhelpful reasons.
Figure 3: Boxplot of tweet post reactions in COMMUNITYNOTES. We measure popularity using the number of replies, likes, and retweets. These posts receive a median of approximately 3k likes, 300 replies, and 700 retweets, indicating that notes are frequently attached to already popular content.
Figure 4: The language distribution of COMMUNITYNOTES: en=English, fr=French, es=Spanish, ja=Japanese, pt=Portuguese, de=German, Other=languages that appear less than 1,000 times.
We present the distribution of helpfulness reasons in COMMUNITYNOTES in Figure˜2. The most frequent helpful reason categories are helpfulClear, helpfulImportantContext and helpfulAddressClaim. They account for the majority of helpful instances in our dataset.
@@ 3.3 Dataset Statistics
Mixed helpful / unhelpful notes 04,557 04.34%
@@
The most frequent not helpful reasons in our dataset are notHelpfulNoteNotNeeded, notHelpfulMissingKeyPoints, and notHelpfulOpinionSpeculationOrBias. They form a distinct middle tier in COMMUNITYNOTES. The remaining categories form a long tail with substantially lower frequencies.
Beyond note level statistics, we also examine the popularity of the posts targeted by notes in Figure˜3. We measure popularity using the number of replies, likes, and retweets. These posts receive a median of approximately 3k likes, 300 replies, and 700 retweets, indicating that notes are frequently attached to already popular content. As shown in Figure˜3, The mean reaction counts are substantial (e.g., $\approx$ 1,200 replies, $\approx$ 2,800 retweets, and $\approx$ 20,000 likes), indicating that posts in the dataset receive great public attention on average. However, the median for all three metrics is near zero. This large gap between the mean and median confirms the existence of a long tail of highly viral tweets. Given that such posts are flagged as potentially misleading, this highlights the importance of providing clear and helpful explanations at scale.
Figure˜4 illustrates the language distribution of COMMUNITYNOTES. English is the dominant language, accounting for 57.3% of all notes, followed by Japanese (10.5%) and Spanish (9.9%). This indicates that our final dataset effectively captures the multilingual nature of crowd-sourced explanations.
@@ 3.4 Benchmarking
To establish a baseline for understanding the model performance in predicting the note helpfulness on COMMUNITYNOTES, we conducted a comprehensive set of experiments using both Small Language Models (SLMs) (BERT-style encoder-only architectures in our paper) and decoder-only Large Language Models (LLMs). This setup provides a clear view of how different model architectures and capacities affect performance on our helpfulness evaluation task.
For SLMs, we formulate the task as multi-task classification, jointly predicting: (1) a binary helpfulness label; and (2) a multi-label reason classification, across 18 predefined categories. Inputs are constructed as a concatenated sequence in the format “[CLS] Claim: {claim} [SEP] Note: {note_text}”. For LLMs, we fine-tune using LoRA fine-tuning (Hu et al., 2022) (see Appendix LABEL:tab:original_prompt for detailed prompts).
We partitioned our experiments by model and language setting. For English, we experimented with SLMs including BERT (Devlin et al., 2019), ModernBERT (Warner et al., 2025), RoBERTa (Liu et al., 2019) and DeBERTa (both base and large) (He et al., 2023), For multilingual evaluation, we used BERT-base-multilingual (mBERT) and XLM-RoBERTa. We also evaluated LLMs including Llama3.1-8B (Dubey et al., 2024) and Mistral-7B (Jiang et al., 2023) for both English and multilingual settings on COMMUNITYNOTES.
Language Model Helpfulness F1 Reason P Reason R Reason F1
@@
Llama3.1-8B 0.928 0.653 0.652 0.653
Mistral-7B 0.926 0.647 0.647 0.647
Table 5: Benchmark results for helpfulness and reasons prediction. Helpfulness means the binary classification of notes helpfulness. Reason means multi-label classification of helpfulness reasons. P is Precision and R is Recall.
@@
As shown in Table˜5, models perform well on helpfulness prediction, with most of them achieving F1 above 0.88. This suggests that this task is relatively easy. In contrast, reason prediction is much more challenging, as reflected by substantially lower F1 for all models, which consistently fall below 0.70.
Comparing SLMs to LLMs: Table˜5 shows that LLMs such as Llama3.1-8B and Mistral-7B outperform smaller SLMs on helpfulness prediction. In paticular, Mistral-7B achieves the highest F1 score (0.920), demonstrating the advantages of larger models. However, LLMs perform worse than SLMs in reason prediction, with F1 remaining comparable to those of the best-performing SLMs such as DeBERTa-large.
Within the SLM families (BERT, RoBERTa, DeBERTa), larger models consistently outperform their base-sized counterparts. For example, RoBERTa-large achieves better scores than RoBERTa-base across all metrics. This highlights the importance of model capacity for capturing the nuances required for this complex task. The table also shows that DeBERTa models generally perform at or near the top among the SLMs, particularly for the reasons task, with the highest F1 Score of 0.665 among all SLMs.
@@
Figure 5: Our framework for automatic reason definition generation and optimization. (1) Seed reason definitions are generated from sampled post-note pairs. (2) The definitions are refined via PromptAgent. (3) The optimized definitions are embedded and fused with note representations for joint helpfulness and reason prediction.